# Rich Human Feedback for Text-to-Image Generation

### 15 Dec 2023 v1

@@ Abstract
Recent Text-to-Image (T2I) generation models such as Stable Diffusion and Imagen have made significant progress in generating high-resolution images based on text descriptions. However, many generated images still suffer from issues such as artifacts/implausibility, misalignment with text descriptions, and low aesthetic quality. Inspired by the success of Reinforcement Learning with Human Feedback (RLHF) for large language models, prior works collected human-provided scores as feedback on generated images and trained a reward model to improve the T2I generation. In this paper, we enrich the feedback signal by (i) marking image regions that are implausible or misaligned with the text, and (ii) annotating which words in the text prompt are misrepresented or missing on the image. We collect such rich human feedback on 18K generated images and train a multimodal transformer to predict the rich feedback automatically. We show that the predicted rich human feedback can be leveraged to improve image generation, for example, by selecting high-quality training data to finetune and improve the generative models, or by creating masks with predicted heatmaps to inpaint the problematic regions. Notably, the improvements generalize to models (Muse) beyond those used to generate the images on which human feedback data were collected (Stable Diffusion variants).
@@ 1 Introduction
    The first Rich Human Feedback dataset (RichHF-18K) on generated images (consisting of fine-grained scores, implausibility(artifact)/misalignment image regions, and misalignment keywords), on 18K Pick-a-Pic images. The RichHF-18K data set will be released soon.
@@ 2 Related works
Most recently, Diffusion Models (DMs) [46, 22, 36, 41] have emerged as the SOTA for Image Generation [13]. DMs are trained to generate images progressively from random noise, with the ability to capture more diversity than GANs and achieve good sample quality [13]. Latent Diffusion Models [41] are a further refinement that performs the diffusion process in a compact latent space for more efficiency.
@@ 3.1 Data collection process
In this section, we discuss our procedure to collect the RichHF-18K dataset, which includes two heatmaps (artifact/implausibility and misalignment), four fine-grained scores (plausibility, alignment, aesthetics, overall), and one text sequence (misaligned keywords).
For each generated image, the annotators are first asked to examine the image and read the text prompt used to generate it. Then, they mark points on the image to indicate the location of any implausibility/artifact or misalignment w.r.t the text prompt. The annotators are told that each marked point has an “effective radius” (1/20 of the image height), which forms an imaginary disk centering at the marked point. In this way, we can use a relatively small amount of points to cover the image regions with flaws. Lastly, annotators label the misaligned keywords and the four types of scores for plausibility, image-text alignment, aesthetic, and overall quality, respectively, on a 5-point Likert scale. Detailed definitions of image implausibility/artifact and misalignment can be found in the supplementary materials.
@@
We created detailed annotation guidelines to instruct the annotators regarding the annotation steps, interactions with the web UI, examples of different types of implausibility, artifacts, and misalignment. All the annotators (27 in total) are trained with the annotation guidelines and calibrated, before they perform the annotation in order to reduce annotation discrepancy and improve quality. Our annotation took around 3,000 rater-hours in total. More details about our data collection process are in the supplementary.
@@ 3.3 RichHF-18K: a dataset of rich human feedback
We select a subset of image-text pairs from the Pick-a-Pic dataset for data annotation. Although our method is general and applicable to any generated images, we choose the majority of our dataset to be photo-realistic images, due to its importance and wider applications. Moreover, we also want to have balanced categories across the images. To ensure balance, we utilized the PaLI visual question answering (VQA) model [7] to extract some basic features from the Pick-a-Pic data samples. Specifically, we asked the following questions for each image-text pair in Pick-a-Pic. 1) Is the image photorealistic? 2) Which category best describes the image? Choose one in ‘human’, ‘animal’, ‘object’, ‘indoor scene’, ‘outdoor scene’. PaLI’s answers to these two questions are generally reliable under our manual inspection. We used the answers to sample a diverse subset from Pick-a-Pic, resulting in 17K image-text pairs. We randomly split the 17K samples into two subsets, a training set with 16K samples and a validation set with 1K samples. The distribution of the attributes of the 16K training samples is shown in Fig. 2. Additionally, we collect rich human feedback on the unique prompts and their corresponding images from the Pick-a-Pic test set as our test set. In total, we collected rich human feedback on the 18K image-text pairs from Pick-a-Pic. Our RichHF-18K dataset consists of 16K training, 1K validation, and 1K test samples.
@@ 3.4 Data statistics of RichHF-18K
To analyze the rating agreement among annotators for an image-text pair, we calculate the maximum difference among the scores: $\text{max}_{\text{diff}}=\max(\text{scores})-\min(\text{scores})$ where scores are the three score labels for an image-text pair. We plot the counts of the $\text{max}_{\text{diff}}$ for each score type in Fig. 4. We can see that around 25% of the samples have perfect annotator agreement and around 85% of the samples have good annotator agreement ($\text{max}_{\text{diff}}$ is less than or equal to 0.25 after the standardization or 1 in the 5-point Likert scale).
@@ 4.1.2 Model variants
Another approach is to use a single head for each prediction type, i.e., three heads for heatmap, score, and misalignment sequence, respectively. To inform the model of the fine-grained heatmap or score type, we augment the prompt with the output type. More specifically, we prepend a task string (e.g., ‘implausibility heatmap’) to the prompt for each particular task of one example and use the corresponding label as the training target. During inference, by augmenting the prompt with the corresponding task string, the single heatmap (score) head can predict different heatmaps (scores). As we show in the experiments, this augmented prompt approach can create task-specific vision feature maps and text encodings, which performs significantly better in some of the tasks.
@@ 4.2 Experiments
Figure 6: Examples of implausibility heatmaps
@@
PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC
@@
CLIP -- -- -- -- 0.185 0.130 -- --
@@
Table 2: Implausibility heatmap prediction results on the test set. $GT=0$ refers to empty implausibility heatmap, i.e., no artifacts/implausibility, for ground truth. $GT>0$ refers to heatmaps with artifacts/implausibility, , for ground truth.
@@
Table 3: Text misalignment heatmap prediction results on the test set. $GT=0$ refers to empty misalignment heatmap, i.e., no misalignment, for ground truth. $GT>0$ refers to heatmaps with misalignment, for ground truth.
@@ 4.2.1 Experimental setup
For comparison, we finetune two ResNet-50 models [18], with multiple fully connected layers and deconvolution heads to predict the scores and heatmaps accordingly. For the misalignment score, we use the CLIP score as a baseline. For misalignment heatmap prediction, we use CLIP gradient [45] map as a baseline.
@@ 5 Learning from rich human feedback
In this section, we investigate whether the predicted rich human feedback (e.g., scores and heatmaps) can be used to improve image generation. To ensure that the gains from our RAHF model generalize across generative model families, we use Muse [6] as our target model to improve, which is based on a masked transformer architecture and thus different from the Stable Diffusion model variants in our RichHF-18K dataset.
@@
In Fig. 9 (a)-(b), we show one example of finetuning Muse with our predicted plausibility score (threshold=0.8). To quantify the gain from Muse finetuning, we used 100 new prompts to generate images, and asked 6 annotators to perform side-by-side comparisons (for plausibility) between two images from the original Muse and the fine-tuned Muse respectively. The annotators choose from five possible responses (image A is significantly/slightly better than image B, about the same, image B is slightly/significantly better than image A), without knowledge of which model is used to generate the image A/B. The results in Tab. 5 demonstrate that the finetuned Muse with RAHF plausibility scores processes significantly fewer artifacts/implausibility than the original Muse.
@@ 9 Experimental details
We train the model on the datasets with a batch size of 256 for 20K iterations. We utilize the AdamW optimizer with a base learning rate of 0.015. We linearly increase the learning rate from 0 to the base learning rate in the first 2000 iterations, and then decrease the learning rate with a reciprocal square root scheduler w.r.t the number of iterations.

### 09 Apr 2024 v2

@@ Abstract
Recent Text-to-Image (T2I) generation models such as Stable Diffusion and Imagen have made significant progress in generating high-resolution images based on text descriptions. However, many generated images still suffer from issues such as artifacts/implausibility, misalignment with text descriptions, and low aesthetic quality. Inspired by the success of Reinforcement Learning with Human Feedback (RLHF) for large language models, prior works collected human-provided scores as feedback on generated images and trained a reward model to improve the T2I generation. In this paper, we enrich the feedback signal by (i) marking image regions that are implausible or misaligned with the text, and (ii) annotating which words in the text prompt are misrepresented or missing on the image. We collect such rich human feedback on 18K generated images (RichHF-18K) and train a multimodal transformer to predict the rich feedback automatically. We show that the predicted rich human feedback can be leveraged to improve image generation, for example, by selecting high-quality training data to finetune and improve the generative models, or by creating masks with predicted heatmaps to inpaint the problematic regions. Notably, the improvements generalize to models (Muse) beyond those used to generate the images on which human feedback data were collected (Stable Diffusion variants). The RichHF-18K data set will be released in our GitHub repository: https://github.com/google-research/google-research/tree/master/richhf_18k.
@@ 1 Introduction
    The first Rich Human Feedback dataset (RichHF-18K) on generated images (consisting of fine-grained scores, implausibility(artifact)/misalignment image regions, and misalignment keywords), on 18K Pick-a-Pic images.
@@ 2 Related works
Most recently, Diffusion Models (DMs) [47, 22, 37, 42] have emerged as the state-of-the-art (SOTA) for Image Generation [13]. DMs are trained to generate images progressively from random noise, with the ability to capture more diversity than GANs and achieve good sample quality [13]. Latent Diffusion Models [42] are a further refinement that performs the diffusion process in a compact latent space for more efficiency.
@@ 3.1 Data collection process
In this section, we discuss our procedure to collect the RichHF-18K dataset, which includes two heatmaps (artifact/implausibility and misalignment), four fine-grained scores (plausibility, alignment, aesthetics, and overall score), and one text sequence (misaligned keywords).
For each generated image, the annotators are first asked to examine the image and read the text prompt used to generate it. Then, they mark points on the image to indicate the location of any implausibility/artifact or misalignment w.r.t the text prompt. The annotators are told that each marked point has an “effective radius” (1/20 of the image height), which forms an imaginary disk centering at the marked point. In this way, we can use a relatively small amount of points to cover the image regions with flaws. Lastly, annotators label the misaligned keywords and the four types of scores for plausibility, image-text alignment, aesthetic, and overall quality, respectively, on a 5-point Likert scale. Detailed definitions of image implausibility/artifact and misalignment can be found in the supplementary materials. We designed a web UI, as shown in Fig. 1, to facilitate data collection. More details about data collection process can be found in the supplementary materials.
@@ 3.3 RichHF-18K: a dataset of rich human feedback
We select a subset of image-text pairs from the Pick-a-Pic dataset for data annotation. Although our method is general and applicable to any generated images, we choose the majority of our dataset to be photo-realistic images, due to its importance and wider applications. Moreover, we also want to have balanced categories across the images. To ensure balance, we utilized the PaLI visual question answering (VQA) model [7] to extract some basic features from the Pick-a-Pic data samples. Specifically, we asked the following questions for each image-text pair in Pick-a-Pic. 1) Is the image photorealistic? 2) Which category best describes the image? Choose one in ‘human’, ‘animal’, ‘object’, ‘indoor scene’, ‘outdoor scene’. PaLI’s answers to these two questions are generally reliable under our manual inspection. We used the answers to sample a diverse subset from Pick-a-Pic, resulting in 17K image-text pairs. We randomly split the 17K samples into two subsets, a training set with 16K samples and a validation set with 1K samples. The distribution of the attributes of the 16K training samples is shown in the supplementary materials. Additionally, we collect rich human feedback on the unique prompts and their corresponding images from the Pick-a-Pic test set as our test set. In total, we collected rich human feedback on the 18K image-text pairs from Pick-a-Pic. Our RichHF-18K dataset consists of 16K training, 1K validation, and 1K test samples.
@@ 3.4 Data statistics of RichHF-18K
To analyze the rating agreement among annotators for an image-text pair, we calculate the maximum difference among the scores: $\text{max}_{\text{diff}}=\max(\text{scores})-\min(\text{scores})$ where scores are the three score labels for an image-text pair. We plot the histogram of $\text{max}_{\text{diff}}$ in Fig. 4. We can see that around 25% of the samples have perfect annotator agreement and around 85% of the samples have good annotator agreement ($\text{max}_{\text{diff}}$ is less than or equal to 0.25 after the standardization or 1 in the 5-point Likert scale).
@@ 4.1.2 Model variants
Another approach is to use a single head for each prediction type, i.e., three heads in total, for heatmap, score, and misalignment sequence, respectively. To inform the model of the fine-grained heatmap or score type, we augment the prompt with the output type. More specifically, we prepend a task string (e.g., ‘implausibility heatmap’) to the prompt for each particular task of one example and use the corresponding label as the training target. During inference, by augmenting the prompt with the corresponding task string, the single heatmap (score) head can predict different heatmaps (scores). As we show in the experiments, this augmented prompt approach can create task-specific vision feature maps and text encodings, which performs significantly better in some of the tasks.
@@ 4.2 Experiments
Figure 5: Examples of implausibility heatmaps. Prompt: photo of a slim asian little girl ballerina with long hair wearing white tights running on a beach from behind nikon D5
@@
PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$ PLCC $\uparrow$ SRCC $\uparrow$
@@
PickScore (off-the-shelf) 0.0098 0.0280 0.131 0.140 0.346 0.340 0.202 0.226
CLIP (off-the-shelf) -- -- -- -- 0.185 0.130 -- --
CLIP (fine-tuned) 0.390 0.378 0.357 0.360 0.398 0.390 0.353 0.352
@@
Table 2: Implausibility heatmap prediction results on the test set. $GT=0$ refers to empty implausibility heatmap, i.e., no artifacts/implausibility (69 out of 995 test samples are empty), for ground truth. $GT>0$ refers to heatmaps with artifacts/implausibility, for ground truth.
@@
Table 3: Text misalignment heatmap prediction results on the test set. $GT=0$ refers to empty misalignment heatmap, i.e., no misalignment (144 out of 995 test samples are empty), for ground truth. $GT>0$ refers to heatmaps with misalignment, for ground truth.
@@ 4.2.1 Experimental setup
For comparison, we finetune two ResNet-50 models [18], with multiple fully connected layers and deconvolution heads to predict the scores and heatmaps, respectively. We also use the off-the-shelf PickScore model [31] to compute the PickScores and calculate the metrics w.r.t each of our four ground truth scores. We use the off-the-shelf CLIP model [40] as a baseline to compute the cosine similarity of the image and text embeddings and use it to calculate the text-image alignment metric, as the CLIP cosine similarity is designed to reflect the alignment between images and prompts. Besides, we also fine-tune a CLIP model to predict the four types of scores using our training dataset. For misalignment heatmap prediction, we use CLIP gradient [46] map as a baseline.
@@ 5 Learning from rich human feedback
In this section, we investigate whether the predicted rich human feedback (e.g., scores and heatmaps) can be used to improve image generation. To ensure that the gains from our RAHF model generalize across generative model families, we mainly use Muse [6] as our target model to improve, which is based on a masked transformer architecture and thus different from the Stable Diffusion model variants in our RichHF-18K dataset.
@@
In Fig. 8 (a)-(b), we show one example of finetuning Muse with our predicted plausibility score (threshold=0.8). To quantify the gain from Muse finetuning, we used 100 new prompts to generate images, and asked 6 annotators to perform side-by-side comparisons (for plausibility) between two images from the original Muse and the fine-tuned Muse respectively. The annotators choose from five possible responses (image A is significantly/slightly better than image B, about the same, image B is slightly/significantly better than image A), without knowledge of which model is used to generate the image A/B. The results in Tab. 5 demonstrate that the finetuned Muse with RAHF plausibility scores possesses significantly fewer artifacts/implausibility than the original Muse.
@@ 8.3 Additional details
Annotation interface
More details
We created detailed annotation guidelines to instruct the annotators regarding the annotation steps, interactions with the web UI, examples of different types of implausibility, artifacts, and misalignment. All the annotators (27 in total) are trained with the annotation guidelines and calibrated, before they perform the annotation in order to reduce annotation discrepancy and improve quality. Our annotation took around 3,000 rater-hours in total. To improve the effectiveness of the collected dataset and control the time spent on annotation, we filter out any image-text pairs that have a text prompt with less than 3 words or more than 20 words. We also filter out non-English prompts or any prompts containing emoji.
@@
Additional details of data collection
The distribution of the attributes of the 16K training samples is shown in Fig. 10. We can see a relatively balanced distribution of the types of content in the generated images in our dataset.
@@ 9 Experimental details
We train the model on the datasets with a batch size of 256 for 20K iterations. We utilize the AdamW optimizer with a base learning rate of 0.015. We linearly increase the learning rate from 0 to the base learning rate in the first 2000 iterations, and then decrease the learning rate with a reciprocal square root scheduler w.r.t the number of iterations. We trained the model using 64 Google Cloud TPU v3 chips.
@@
Finetuning generative models with predicted scores
To generate the training prompt set, we provide five hand-crafted seed prompts as examples and then ask PaLM 2 [1] to generate similar textual prompts. We include additional instructions that specify the prompt length and the object category. We then explain why we do not use existing benchmark datasets for training. Theoretically, we can get an infinite number of prompts using the prompt synthesis technique we proposed above. Existing datasets are 1) relatively small (e.g., TIFA [24] has 4k prompts, Davidsonian Scene Graph (DSG) [10] has only 1k prompts), or 2) containing prompts that are simple and not diverse enough, for example, only measuring single objects in Parti benchmark [59]. This motivates us to synthesize a larger set of diverse prompts for training purposes. For the 100 prompts for our human evaluation, they are sampled from the existing benchmark: TIFA [24]. We only did our evaluation on 100 prompts due to the high cost of the human annotation.